OpenStudy (anonymous):

-4k-11=9k+2 Help me?

OpenStudy (lilg132):

dont forget to click good answer and let me know if you dont understand

OpenStudy (turingtest):

c'mon lilg, you made a pretty big mistake...

OpenStudy (turingtest):

@Bee Lovee the answer is wrong above

OpenStudy (lilg132):

oops wait let me check

OpenStudy (lilg132):

my baddddddddd let me start again

OpenStudy (turingtest):

that's what you get for trying to answer 10 questions at once lol!

OpenStudy (lilg132):

-4k-11=9k+2 -11 - 2 = 9k + 4k -13 = 13k -13 / 13 = 13k / 13 -1 = k correct answer turing?

OpenStudy (turingtest):

yessir!

OpenStudy (lilg132):

lol thanks for spotting my mistake :)

OpenStudy (turingtest):

thanks for correcting it :)